I need help please! I don't understand this. The problem is:
7x+2y= 10
-7x+y= -16

Answer by mananth(16949)

7 x + 2 y = 10 .............1
-7 x + 1 y = -16 .............2
Eliminate y
multiply (1)by 1
Multiply (2) by -2
7x+2y=10
14x-2y= 32
Add the two equations
21x=42
/21
x=2
plug value of x in (1)
7 x + 2 y = 10
14 + 2 y = 10
2 y = 10-14
2 y = -4
y = -2
m.ananth@hotmail.ca